40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Apr 25 03:14:46 2014, ericq, Update, LSC, locking activity 
    Reply  Fri Apr 25 13:18:08 2014, Dark Jamie, Update, LSC, locking activity m10008_f1_bg.jpg
Message ID: 9855     Entry time: Fri Apr 25 13:18:08 2014     In reply to: 9853
Author: Dark Jamie 
Type: Update 
Category: LSC 
Subject: locking activity 

Quote:

[ericq, Jenne, Zach]

We spent some time tonight trying to push our CARM locking further, to little avail. DARM/CARM loop oscillations kept sneaking up on us. We measured some MC2 motion -> REFL11 Transfer Functions to see if we could see CARM plant features; plots will come in the near future...

 Probably things would have worked better if you would have gotten your hair done at the same place as me.

Attachment 1: m10008_f1_bg.jpg  48 kB  | Hide | Hide all
m10008_f1_bg.jpg
ELOG V3.1.3-